OpenStudy (anonymous):

write the slope-intercept form of the line through the points (-7,1) and (-4, -2)

OpenStudy (anonymous):

okay, so for this, the first thing you would want to do is find the slope. the formula for that would be \[\frac{ y2-y1 }{ x2-x1 }\] when you plug the points in, you would get -1 as your slope. then you would have to find the intercepts by using the point-slope formula. \[y-y1=m(x-x1)\] then, you can easily put it into slope-intercept form :) let me know if you're still having trouble with anything

OpenStudy (anonymous):

so how would the answer look like if u plugged in back the #s

OpenStudy (anonymous):

\[\frac{ -2-1 }{ -4-(-7) }\] ^^thats your slope, which is equal to -1 then to plug it into the other formula, you would choose one of the points (I chose the first) and it would look like this: \[y-(1)=-1(x-(-7))\] if you solve that out, you should get y=-x-6

OpenStudy (anonymous):

ty